### v1

@@ 5.3 Ablation Studies
EoE 100 95.3 95.7 94.4 92.5 91.6 88.3 87.5 86.8 86.0
WAVE-CRE 100 96.8 94.5 93.3 91.0 89.5 88.6 87.6 86.4 85.4
WAVE++ 100 97.9 96.1 95.1 94.0 92.7 91.7 90.8 89.6 88.3
@@
EoE 100 95.2 87.5 85.0 82.0 82.4 83.3 82.7 84.1 82.4
WAVE-CRE 100 96.8 91.1 86.8 84.9 83.6 82.1 80.4 80.2 79.2
WAVE++ 100 95.9 92.0 90.5 88.7 87.9 86.9 85.9 85.2 84.8
Task Identity Inference
Instead of employing a conventional MLP head for task classification, we utilize a cascade voting mechanism for task identity prediction. This approach not only reduces training overhead but also enhances task identification by aggregating predictions from all prompt pools during the selection process. To assess its effectiveness relative to an MLP-based method, such as WAVE-CRE (Le et al., 2025a), we conducted a controlled experiment measuring task classification accuracy, with results presented in Table 8.
The results demonstrate that WAVE++ significantly outperforms WAVE-CRE, with an approximate improvement of 3% on the FewRel dataset and 4% on TACRED after training on all tasks. In addition to these gains in raw accuracy, the cascade voting mechanism exhibits greater stability compared to the MLP-based classifier. In particular, the MLP-based approach shows occasional sharp declines in accuracy, indicative of its sensitivity to distribution shifts when new tasks are introduced. In contrast, the voting mechanism maintains consistent performance, highlighting its robustness in managing heterogeneous task distributions. Moreover, WAVE++ surpasses EoE in terms of task prediction accuracy. These findings underscore the effectiveness and resilience of the cascade voting mechanism in complex, multi-task scenarios, where adaptability and robustness to distribution shifts are crucial.
6 Conclusion
In this paper, we introduce a novel perspective on prompt-based continual learning by interpreting existing methods through the lens of mixture of experts models. We show that contemporary prompt-based continual learning approaches can be viewed as a special case of MoE architectures. Building on this insight, we present WAVE++, which shares key conceptual similarities with sparse MoE designs. Our framework addresses several challenges in prompt-based continual learning, including suboptimal task identification, sensitivity to distribution shifts, and handling within-task diversity. Specifically, we mitigate these issues by employing a cascade voting mechanism for task prediction and using task-specific prompt pools, complemented by label descriptions to enhance continual relation extraction performance.
Despite these advancements, there remain hurdles to improving the efficiency and effectiveness of WAVE++. Notably, the method heavily depends on the design of prompt pools and experts, suggesting opportunities to refine and optimize these components. While our approach attenuates catastrophic forgetting, retaining knowledge of previous tasks remains challenging, echoing issues highlighted in prior studies. In particular, if prompt pools are not properly applied during testing, model forgetfulness may reemerge. Moreover, although prompt pools enhance the expressiveness of our method, the current prefix-tuning experts are relatively simplistic. Future work may explore more sophisticated expert architectures to further strengthen the model.

### v2

@@ Appendix F Performance under Incorrect Task Prediction
Percentage 0 18.2 10.0 13.8 11.8 9.4 11.8 11.9 12.4 14.4
Correct final prediction cases 0 4 4 8 8 8 13 14 16 20
WAVE-CRE Errors in task prediction 0 22 40 58 68 85 110 118 129 139
Percentage 0 39.3 22.0 17.1 18.2 17.3 18.4 20.5 15.3 19.2
Correct final prediction cases 0 7 8 8 11 15 18 22 18 25
WAVE++ Errors in task prediction 0 18 37 47 60 89 102 110 119 133
@@
Percentage 0 14.3 5.0 15.0 10.7 7.7 8.0 7.7 8.3 12.0
Correct final prediction cases 0 1 1 3 3 2 2 2 2 3
WAVE-CRE Errors in task prediction 0 7 20 20 28 26 25 26 24 25
Percentage 0 25.0 25.0 26.3 19.0 14.3 16.7 16.7 23.5 20.0
Correct final prediction cases 0 2 5 5 4 3 3 3 4 3
WAVE++ Errors in task prediction 0 8 20 19 21 21 18 18 17 15
Table 11: Final accuracy under incorrect task prediction
We present detailed results on the model’s performance when the initial task prediction is incorrect. Table 11 reports three key metrics: (i) the number of task prediction errors, (ii) the number of correct final relation predictions despite an incorrect task prediction, and (iii) the percentage of correct predictions among all cases of incorrect task prediction.
Across both datasets, WAVE++ consistently yields fewer task prediction errors than WAVE-CRE, highlighting the advantage of cascade voting over the MLP classifier. Moreover, when the task prediction is incorrect, WAVE++ achieves substantially higher final classification accuracy, outperforming WAVE-CRE by 3–10% on FewRel and 7–20% on TACRED. These results demonstrate the robustness of WAVE++ with cascade voting, even in the presence of task prediction errors.
Appendix G Statistical Significance Tests
To assess whether the proposed method WAVE++ significantly outperforms prior approaches, we conducted paired t-tests against two baselines, WAVE-CRE and HiDe-Prompt, on the TACRED and FewRel datasets. Each comparison was performed over five independent runs with different random seeds.
For each test, the null hypothesis $H_{0}$ states that the mean performance difference between WAVE++ and the corresponding baseline is zero, while the alternative hypothesis $H_{1}$ posits that WAVE++ outperforms the baseline. We adopt a significance level of $\alpha=0.05$. As reported in Table 12, all obtained p-values are substantially below the significance threshold, demonstrating that the performance gains of WAVE++ over both WAVE-CRE and HiDe-Prompt are statistically significant.
Dataset TACRED FewRel
Method WAVE-CRE HiDe-Prompt EoE WAVE-CRE HiDe-Prompt EoE
pp-value 3.6e-04 9.5e-04 5.4e-03 2.8e-06 5.2e-04 1.5e-04
Table 12: Paired t-test results evaluating whether WAVE++ significantly outperforms on the TACRED and FewRel datasets.
Appendix H Time Analysis
Method Training Time (h) Inference Latency (ms)
TACRED FewRel TACRED FewRel
L2P 0.9 6.1 26.6 27.2
EoE 1.2 6.5 39.5 40.6
WAVE-CRE 2.5 7.4 28.7 29.8
WAVE++ 1.4 6.8 40.5 41.2
Table 13: Training time and inference latency on the TACRED and FewRel datasets.
In this section, we evaluate the computational efficiency of the proposed method by comparing training time and inference latency, as reported in Table 13. As shown, WAVE++ achieves training time and inference latency comparable to the state-of-the-art method EoE. Moreover, both EoE and WAVE++ require substantially less training time than WAVE-CRE, as they eliminate the need to train a task predictor during the learning phase. This efficiency gain, however, comes at the cost of increased inference latency, since cascade voting is employed at test time. This trade-off between training efficiency and inference overhead is shared by both EoE and WAVE++.